mathslover (mathslover):

Find the least value of the expression \(2\log_{10} x - \log_x {0.01} \) for x >1

mathslover (mathslover):

@hartnn @Miracrown

mathslover (mathslover):

For this question, \(2\log_{10} x + 2\log_x {10} \) oh..

hartnn (hartnn):

that log has base 10, right ?

mathslover (mathslover):

\(2\cfrac{1}{\log_x {10} } + 2\log _x {10}\) \(2 \left( \cfrac{1 + (\log_x {10})^2}{\log_x {10}}\right)\)

mathslover (mathslover):

Yep.. its 10.

mathslover (mathslover):

I don't : whether I am on right track here , if yes, then ho wto continue?

hartnn (hartnn):

and you cannot use calculus ?

mathslover (mathslover):

Oh... maxima minima?

hartnn (hartnn):

least value means minima

mathslover (mathslover):

Okay right.. so, if I differentiate it and equate to zero

hartnn (hartnn):

if you can use calculus

hartnn (hartnn):

u = log x 2(u+1/u) diff. w.r.t u equate to 0

mathslover (mathslover):

Oh.. k i will try now.

mathslover (mathslover):

\(\cfrac{d}{du} 2(u + \cfrac{1}{u}) \) = \( 2( 1 + (-1) \cfrac{1}{u^2})\) = 0 1 = 1/u^2 u^2 = 1

mathslover (mathslover):

2(u + 1/u ) = 2 (1+1/1) = 2(2) = 4

mathslover (mathslover):

2(1-1/u^2) -> differentiating it again and putting u = 1 we get : 4 / u^3 put u = 1 so, 4/1 = positive so, u = 1 is the minima , right @hartnn ?

mathslover (mathslover):

O.O I just got to know another soln for this : (courtsey @vishweshshrimali5 ) \((u + \cfrac{1}{u} ) \ge 2 \) So, 2(u+1/u) \(\ge\) 4 thus, 4 was the answer...

mathslover (mathslover):

o.O it was much easier with the second method... but, yep, I always like calculus methods @hartnn ;)

hartnn (hartnn):

nice!